# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 8. Quasi-convex functions and subharmonics
    Let $u\in\mathcal{F}(X)$, $v\in\mathcal{G}(X)$. Since the result is local, it is sufficient to show that, for each $x\in X$, one has $u+v\in\mathcal{H}(U)$, where $U$ is any neighborhood of $x$. Hence by upper semicontinuity we may assume that $u$ and $v$ are bounded above. Also, we may assume that $u$ and $v$ are bounded below, since it suffices to consider, for any positive integer $m$, the subharmonics $u\vee(\varphi-m)\vcentcolon=\max\{u,\varphi-m\}\in\mathcal{F}(U)$ and $v\vee(\psi-m)\vcentcolon=\max\{v,\psi-m\}\in\mathcal{G}(U)$, for some bounded quadratic functions $\varphi\in\mathcal{F}(U)$ and $\psi\in\mathcal{G}(U)$, and then take the limit $m\to\infty$. We have already highlighted in Remark 5.11 that the existence of such quadratic functions is guaranteed by the topological property (T) and that their translations by negative constants are still subsolutions by the sliding property yielded by the properness condition (N), while the truncated functions are subharmonic thanks to the maximum property of constant coefficient subequation. With these assumptions, by Proposition 8.8 there exist sequences $\{u_{j}\}$, $\{v_{j}\}$ of quasi-convex subharmonics near $x$ that converge monotonically downward to $u,v$ respectively, as $j\to\infty$. By Theorem 8.3 we know that $u_{j}+v_{j}\in\mathcal{H}(X)$ and by the decreasing sequence property we conclude $u+v\in\mathcal{H}(X)$. ∎
    We conclude this section with some remarks on the utility, limitations and generalizations of the subharmonic addition theorem for constant coefficient subequations in Corollary 8.9. First, we give one important use of Corollary 8.9, which concerns the valiidty of the comparison principle and will be the subject of Section 9.
@@
    If a subequation constraint set $\mathcal{F}$ on $X$ (not necessarily with constant coefficients) is $\mathcal{M}$-monotone for some constant coefficient monotonicity cone subequation $\mathcal{M}$ as in (6.2); that is
    $\mathcal{F}_{x}+\mathcal{M}\subset\mathcal{F}_{x},\quad\forall\,x\in X,$
    then duality gives the jet addition formula
    $\mathcal{F}_{x}+\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{M}},\quad\forall\,x\in X.$
@@
    $\mathcal{F}(X)+\widetilde{\mathcal{F}}(X)\subset\widetilde{\mathcal{M}}(X).$ (8.2)
    Again using duality, the validity of the comparison principle on $\Omega\Subset X$ can be reduced to the validity of the Zero Maximum Principle (ZMP) for $z\in\operatorname{USC}(\overline{\Omega}\cap\widetilde{\mathcal{M}}(\Omega)$:
    $z\leqslant 0\ \text{on $\partial\Omega$}\ \ \Rightarrow\ \ z\leqslant 0\ \text{on $\partial\Omega$},$ (8.3)
    by considering every $z:=u+v$ with $u\in\operatorname{USC}(\overline{\Omega})\cap\mathcal{F}(\Omega)$ and $v\in\operatorname{USC}(\overline{\Omega})\cap\widetilde{\mathcal{F}}(\Omega)$. Corollary 8.9 says that if $\mathcal{F}$ has constant coefficients, then (8.2) indeed holds. This is the monotonicty-duality method for proving comparison, which is extensively treated in the constant coefficient setting in [8] including a classification of monotonicity cone subequations $\mathcal{M}$ and the validity of the (ZMP) in (8.3) for the dual cone subharmonics.
@@
    In the constant coefficient setting, the quasi-convex approximation preserves $\mathcal{F}$-subharmonicity. This is essentually due to validity of the translation property of Proposition 7.4 (viii), which one loses when passing to nonconstant coefficients (as seen at the end of the proof of this proposition). In the variable coefficient setting, one can extend the quasi-convex approximation technique to obtain the subharmonic addition theorem and its consequences provided that the variable coefficient subequation $\mathcal{F}$ is both $\mathcal{M}$-monotone (for some monotonicty cone subequation $\mathcal{M}$) and fiberegular.
    The $\mathcal{M}$-monotonicty of $\mathcal{F}$ always holds for variable coefficient pure second order and gradient-free subequations, but when there is gradient dependence, one also needs directionality, as studied in [11]. Moreover, in this case, the assupmtion of $\mathcal{M}$-monotonicity is also needed for the monotonicity-duality method for proving the comparison principle, as noted in Remark 8.10.
    The fiberegularity of $\mathcal{F}$ is just the continuity of the fiber map $\Theta:X\to\wp(\mathcal{J}^{2})$ defined by $\Theta_{x}:=\mathcal{F}_{x}$ for each $x\in X$. Using this regularity property and the $\mathcal{M}$-monotonicity, one can replace the translation property by the uniform translation property, which roughly speaking means that if $u\in\mathcal{F}(\Omega)$, then there are small $C^{2}$ strictly $\mathcal{F}$-subharmonic perturbations of all small translates of $u$ which belong to $\mathcal{F}(\Omega_{\delta})$, where $\Omega_{\delta}:=\{x\in\Omega:d(x,\partial\Omega)>\delta\}$ (see Theorem 3.3 of [11] for a precise statement and the proof). This uniform translation property was developed first in the variable coefficient pure second order case in [9] and generalized to the variable gradient-free case in [10], where we stress again that the needed $\mathcal{M}$ monotonicity holds in these two gradient-free cases.
    9. Comparison principles
    In this section, we illustrate some aspects of the monotonicity-duality method for proving the comparison principle in nonlinear potential theories. We gave a synthetic sketch of the method in Remark 8.10 as an application of the Subharmonic Addition Theorem. Here, we will derive the method from first principles. Then we will provide various illustrations of its applications, as described in the introduction to Part II.
@@ 9. Comparison principles
    $\begin{cases}w\in\mathcal{F}(X)+\widetilde{\mathcal{F}}(X)\\ w\leqslant 0\ \text{on $\partial\Omega$}\end{cases}\implies w\leqslant 0\ \text{on $\Omega$},$ (ZMP)
    which merely exploits the fact that $v$ is $\mathcal{F}$-superharmonic if and only if $-v\in\widetilde{\mathcal{F}}(X)$. This suggests that a key role in the proof of comparison could be played by a subharmonic addition theorem; that is, one would like to know if
    $\mathcal{F}_{x}+\widetilde{\mathcal{F}_{x}}\subset\mathcal{G}_{x},\ \ \forall x\in X\implies\mathcal{F}(X)+\widetilde{\mathcal{F}}(X)\subset\mathcal{G}(X).$
@@
    In order to find a suitable $\mathcal{G}$, note that if $\mathcal{F}$ is $\mathcal{M}$-monotone; that is,
    $\mathcal{F}_{x}+\mathcal{M}_{x}\subset\mathcal{F}_{x},\ \ \forall\,x\in X,$ (9.1)
    then by properties (1), (2) and (3) of Proposition 6.2 one has, for each $J\in\mathcal{F}_{x}$,
    $J+\mathcal{M}_{x}\subset\mathcal{F}_{x}\implies\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{M}}_{x}-J,$
    hence
    $\mathcal{F}_{x}+\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{M}}_{x},\ \ \forall\,x\in X$
    and one may choose $\mathcal{G}=\widetilde{\mathcal{M}}$. Moreover, one knows that such a subset $\mathcal{M}$ always exists. Indeed, regardless of the subequation $\mathcal{F}$, thanks to the conditions (P) and (N) a set (with constant fibers) that satisfies (9.1) is
@@
    where
    $\mathcal{N}\vcentcolon=\{r\in{\mathbb{R}}:\ r\leqslant 0\}\quad\text{and}\quad\mathcal{P}\vcentcolon=\{A\in\mathcal{S}(n):\ A\geqslant 0\}.$ (9.2)
    Note that $\mathcal{M}$ is a constant coefficient primitive subequation (it is closed and satisfies property (P)) and $\mathcal{M}$ satisfies the negativity property (N). However, $\mathcal{M}$ it is not a subequation because, having empty interior due to the factor $\{0\}$, it fails to satisfy the topological property (T). This trivial factor says that there may be monotonicity of $\mathcal{F}$ with respect to the gradient variable $p\in{\mathbb{R}}^{n}$.
    Assuming that $\mathcal{M}$ can be enlarged to be a monotonicity cone subequation satisfying (9.1), the only missing piece would be to prove that comparison holds for functions in ~$\widetilde{\mathcal{M}}(X)$. Notice that choosing $\mathcal{M}$ as large as possible facilitates the proof of the ZMP for ~$\widetilde{\mathcal{M}}(X)$, which becomes smaller as $\mathcal{M}$ increases. This represents the main observation to the monotonicity–duality method of Harvey and Lawson, as discussed in [8].
    Having derived the monotonicity-duality method, we will present a few illustrative applications of it. First, a pair of constant coefficient situations coming from [23, 8] for which comparison holds on all bounded domains and the needed (ZMP) for the $\widetilde{\mathcal{M}}$-subharmonics is accomplished by characterizing them. Then, beginning in the next subsection, we will discuss strict comparison for quasi-convex functions in the variable coefficient setting and for semicontinuous functions in the constant coefficient setting.
@@ 9.1. Comparison with constant coefficients and sufficient monotonicity
    We will present two basic examples of classes of $\mathcal{M}$-monotone subequations for which comparison holds on all bounded domains. The classes are determined by the two most elementary monotonicity cone subequations $\mathcal{M}$, which have interesting characterizations of the dual subharmonics ~$\widetilde{\mathcal{M}}(X)$, as known from [23, 8]. These characterizations easily lead to a direct proof that the $\widetilde{\mathcal{M}}$-subharmonics satisfy the (ZMP) on all bounded domains, and hence the comparison principle in all $\mathcal{M}$-monotone potential theories. The two classes discussed here can be included in a much more general result that will be recalled in Remark 9.14 below.
@@
    We have used Remark 5.15 to identify a constant coefficient subequation $\mathcal{F}\subset\mathcal{J}^{2}({\mathbb{R}}^{n})$ with a subset ×${\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$. Notice that such an $\mathcal{F}$ is indeed a subequation because the negativity property (N) is automatic and the topological property (T) follows from the $\mathcal{P}$-monotonicity, which implies the following monotonicity property
    +(×$\mathcal{F}+({\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{P})\subset\mathcal{F},$
    where the property +(×$\mathcal{F}+({\mathbb{R}}\times{\mathbb{R}}^{n}\times\{0\})\subset\mathcal{F}$ is an elegant way to say that $\mathcal{F}$ is of pure second order.
    In addition, since the only component of a jet that is actually relevant to define a subsolution is the one living in $\mathcal{S}(n)$, we will often identify $\mathcal{F}$ with $\mathcal{A}\subset\mathcal{S}(n)$ that satisfies the positivity condition (9.3). Hence, we can say that a pure second-order constant coefficient subequation is a closed subset $\mathcal{A}\subset\mathcal{S}(n)$ which is $\mathcal{P}$-monotone.
    Given $A\in\mathcal{S}(n)$ we denote by $\lambda_{1}(A)$ and $\lambda_{n}(A)$ its minimum and its maximum eigenvalue, respectively (that is, we order the eigenvalues $\lambda_{1}\leqslant\lambda_{2}\leqslant\cdots\leqslant\lambda_{n}$). The most basic of all examples is the convex subequation (cf. (9.2))
    $\mathcal{P}\vcentcolon=\{A\in\mathcal{S}(n):\ \lambda_{1}(A)\geqslant 0\},$
    whose dual subequation is the co-convex or subaffine subequation
    $\widetilde{\mathcal{P}}\vcentcolon=\{A\in\mathcal{S}(n):\ \lambda_{n}(A)\geqslant 0\}.$
    The name subaffine comes from the fact, proved in [23, Proposition 4.5], that an upper semicontinuous function on $X$ is a subsolution of $\widetilde{\mathcal{P}}$ if and only if it is subaffine on $X$; that is
    $\widetilde{\mathcal{P}}(X)=\mathrm{SA}(X),$ (9.4)
    where $\mathrm{SA}(X)$ is the space of all subaffine functions on $X$, in the sense of the following definition.
@@
    $w\leqslant a\ \text{on $\partial\Omega$}\implies w\leqslant a\ \text{on $\Omega$}.$
@@
    We wrote the space of all affine functions as $\mathrm{Aff}$, without specifying their domain because it is known that any $a$ which is affine on an open subset $\Omega$ has a unique extension to the whole space ${\mathbb{R}}^{n}$ (see, e.g., [48]).
    Note that $\mathcal{P}$ is a monotonicity cone subequation for all pure second-order constant coefficient subequations (and it is minimal in this sense), hence a straightforward application of the Subharmonic Addition Theorem (Corollary 8.9) yields the following result. For a proof that uses Słodkowski’s Largest Eigenvalue Theorem and quasi-convex approximation, the reader can instead refer to [23].
    Theorem 9.5 (Subaffine Theorem; [23, Theorem 6.5]).
    If $\mathcal{F}$ is a pure second-order constant coefficient subequation, then for any open set $X\subset{\mathbb{R}}^{n}$
@@
    Let $\Omega\subset{\mathbb{R}}^{n}$ a domain and $w\in\operatorname{USC}(\overline{\Omega})\cap\mathrm{SA}(\Omega)$. Then
    $w\leqslant a\ \text{on $\partial\Omega$}\implies w\leqslant a\ \text{on $\Omega$}$
    for all $a\in\mathrm{Aff}$.
@@
    Assume without loss of generality that $a=0$. We consider an exhaustion of $\Omega$ by an increasing sequence of compact sets $\{K_{j}\}_{j\in\mathbb{N}}$ and set $U_{\delta}\vcentcolon=\{x\in\overline{\Omega}:\ w(x)<\sup_{\partial\Omega}w+\delta\}$, for any $\delta>0$. Since $w\in\operatorname{USC}(\overline{\Omega})$, we know that $U_{\delta}$ is an open neighborhood of $\partial\Omega$ in $\overline{\Omega}$. Therefore $\partial K_{j}\subset U_{\delta}$ for all $j$ sufficiently large and, since $w$ is subaffine, $\sup_{K_{j}}\leqslant\sup_{\partial\Omega}w+\delta$. This proves that $\sup_{\Omega}w\leqslant\sup_{\partial\Omega}w+\delta$; since $\delta$ is arbitrary, $w\leqslant 0$ in $\partial\Omega$ and $w$ is continuous, we conclude that $w\leqslant 0$ on $\overline{\Omega}$, which is the desired conclusion. ∎
    Theorem 9.7 (Comparison for pure second-order subequations).
@@
    $\mathcal{R}+\mathcal{N}\subset\mathcal{R}\quad\text{and}\quad\mathcal{A}+\mathcal{P}\subset\mathcal{A},$ (9.5)
@@
    Just as in Remark 9.2, we have identified the constant coefficient $\mathcal{F}$ with a subset of $\mathcal{J}^{2}$ and it is indeed a subequation, since the remaining topological property (T) again follows by the $(\mathcal{N}\times{\mathbb{R}}^{n}\times\mathcal{P})$-monotonicity of $\mathcal{F}$.
    Moreover, one can again reduce the subequation by suppressing the “silent” gradient variable, which then allows one to further identift $\mathcal{F}$ with a subset of ×${\mathbb{R}}\times\mathcal{S}(n)$. Hence, one can say that a gradient-free constant coefficient subequation is a closed subset ×$\mathcal{F}\subset{\mathbb{R}}\times\mathcal{S}(n)$ which is $\mathcal{Q}$-monotone where
    $\mathcal{Q}:=\mathcal{N}\times\mathcal{P}\subset\mathcal{{\mathbb{R}}}\times\mathcal{S}(n)$ (9.6)
@@
    $\widetilde{\mathcal{Q}}=\{(r,A)\in{\mathbb{R}}\times\mathcal{S}(n):\ r\leqslant 0\ \ \text{or}\ \ A\geqslant 0\}.$ (9.7)
    Also in this case the name comes from a characterization of dual subharmonics. By [8, Theorem 10.7], an upper semicontinuous function on $X$ is $\widetilde{\mathcal{Q}}$-subharmonic if and only if it is subaffine-plus on $X$; that is,
@@
    where $\mathrm{SA}^{+}(X)$ is the space of all subaffine-plus functions on $X$, which are defined as follows.
@@
    the space of affine-plus functions on $K$.
@@
    A function $w\in\operatorname{USC}(X)$ is called subaffine-plus on $X$, and we write $w\in\mathrm{SA}^{+}(X)$, if, for every open subset $\Omega\Subset X$,
    $w\leqslant a_{+}\ \text{on $\partial\Omega$}\implies w\leqslant a_{+}\ \text{on $\Omega$}\qquad\forall a_{+}\!\in\mathrm{Aff}^{+}(\overline{\Omega}).$
@@
    Theorem 9.12 (Subaffine-Plus Theorem; [8, Theorem 10.8]).

### v3

@@ 8.2. The Subharmonic Addition Theorem
Let $u\in\mathcal{F}(X)$, $v\in\mathcal{G}(X)$. Since the result is local, it is sufficient to show that, for each $x\in X$, one has $u+v\in\mathcal{H}(U)$, where $U$ is any neighborhood of $x$. Hence by upper semicontinuity we may assume that $u$ and $v$ are bounded above. Also, we may assume that $u$ and $v$ are bounded below, since it suffices to consider, for any positive integer $m$, the subharmonics $\max\{u,\varphi-m\}\in\mathcal{F}(U)$ and $\max\{v,\psi-m\}\in\mathcal{G}(U)$, for some bounded quadratic functions $\varphi\in\mathcal{F}(U)$ and $\psi\in\mathcal{G}(U)$, and then take the limit $m\to\infty$. Recall that we have already highlighted in Remark 5.11 that the existence of such quadratic functions is guaranteed by the topological property (T), and that their translations by negative constants are still subsolutions by the sliding property (Proposition 7.5(iv)) yielded by the properness condition (N), while the truncated functions are subharmonic thanks to the maximum property (Proposition 7.5(ii)). With these assumptions, by Proposition 8.8 there exist sequences $\{u_{j}\}_{j\in\mathbb{N}}$ and $\{v_{j}\}_{j\in\mathbb{N}}$ of semiconvex subharmonics near $x$ that converge monotonically downward to $u$ and $v$, respectively. By Theorem 8.3 we know that $u_{j}+v_{j}\in\mathcal{H}(X)$ and by the decreasing sequence property (Proposition 7.5(v)) we conclude that $u+v\in\mathcal{H}(X)$. ∎
We conclude this section with some remarks on the utility, limitations and generalizations of the subharmonic addition theorem for constant coefficient subequations in Corollary 8.9. First, we give one important use of it, which concerns the validity of the comparison principle and will be the subject of Chapter 9.
@@
If a subequation constraint set $\mathcal{F}$ on $X$ (not necessarily with constant coefficients) is $\mathcal{M}$-monotone for some constant coefficient monotonicity cone subequation $\mathcal{M}$ as in (6.2), then duality gives the jet addition formula
$\mathcal{F}_{x}+\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{M}}\quad\forall\,x\in X;$
indeed, as for any $J\in\mathcal{F}_{x}$ we have $J+\mathcal{M}\subset\mathcal{F}_{x}$, taking the duals and recalling the properties (1) and (3) of Proposition 6.2, we deduce that $J+\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{M}}$.
@@
$\mathcal{F}(X)+\widetilde{\mathcal{F}}(X)\subset\widetilde{\mathcal{M}}(X),$ (8.2)
and then the validity of the comparison principle on $\Omega\Subset X$ can be reduced to the validity of the Zero Maximum Principle (ZMP) for $z\in\operatorname{USC}(\overline{\Omega})\cap\widetilde{\mathcal{M}}(\Omega)$.
Corollary 8.9 says that if $\mathcal{F}$ has constant coefficients, then (8.2) indeed holds. This is the monotonicty-duality method for proving comparison, which is extensively treated in the constant coefficient setting in [9] including a classification of monotonicity cone subequations $\mathcal{M}$ and the validity of (ZMP) for the dual cone subharmonics.
@@
[…]
The $\mathcal{M}$-monotonicty of $\mathcal{F}$ always holds for variable coefficient pure second order and gradient-free subequations, but when there is gradient dependence, one also needs directionality, as studied in [12]. Moreover, in this case, the assumption of $\mathcal{M}$-monotonicity is also needed for the monotonicity-duality method for proving the comparison principle, as noted in Remark 8.10.
[…]
Chapter 9 Comparison principles
This chapter is dedicated to the proof of comparison principles in general potential theories; that is, the proof that if a subharmonic-superharmonic pair for a subequation $\mathcal{F}$ are ordered on the boundary of a bounded domain, then they are ordered in the interior of the domain. Two distinct situations will be presented. The first situation concerns the use of the monotonicity-duality method which works when $\mathcal{F}$ has sufficient monotoncity. The second situation concerns the extreme case when $\mathcal{F}$ has only the minimal monotonicity that any subequation must have. In this situation, a result on strict comparison is presented for semiconvex subharmonics and dual subharmonics.
9.1. Preliminaries
@@
$\begin{cases}w\in\mathcal{F}(X)+\widetilde{\mathcal{F}}(X)\\ w\leqslant 0\ \text{on $\partial\Omega$}\end{cases}\implies w\leqslant 0\ \text{on $\Omega$},$ (9.1)
[…]
$\mathcal{F}_{x}+\widetilde{\mathcal{F}}_{x}\subset\mathcal{G}_{x}\quad\forall x\in X\quad\implies\quad\mathcal{F}(X)+\widetilde{\mathcal{F}}(X)\subset\mathcal{G}(X).$
@@
In order to find a suitable $\mathcal{G}$, recall that we noted that
$\mathcal{F}_{x}+\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{M}}_{x}\quad\forall\,x\in X$
whenever $\mathcal{F}$ is $\mathcal{M}$-monotone in the sense of (6.2). Therefore, in this case one may choose $\mathcal{G}=\widetilde{\mathcal{M}}$.
Moreover, one knows that such a subset $\mathcal{M}$ always exists. Indeed, regardless of the subequation $\mathcal{F}$, thanks to the conditions (P) and (N) a set (with constant fibers) for which 6.2 holds is
@@
with $\mathcal{N}$ and $\mathcal{P}$ defined as in (6.4); that is, (6.2) always holds for $\mathcal{M}$ being the minimal monotonicity cone $\mathcal{M}_{0}$ defined in (6.3). Note that $\mathcal{M}_{0}$ is a constant coefficient primitive subequation (it is closed and satisfies property (P)) and it also satisfies the negativity property (N); however, we already noted that it is not a subequation because, having empty interior due to the factor $\{0\}$, it fails to satisfy the topological property (T).
Assuming that $\mathcal{M}$ can be enlarged to be a monotonicity cone subequation for which (6.2) holds, the only missing piece would be to prove that the zero maximum principle holds for functions in ~$\widetilde{\mathcal{M}}(X)$. Notice that choosing $\mathcal{M}$ as large as possible facilitates the proof of (ZMP) for ~$\widetilde{\mathcal{M}}(X)$, which becomes smaller as $\mathcal{M}$ increases. This represents the main observation of the monotonicity-duality method of Harvey and Lawson, as discussed in [9].
Having derived the monotonicity-duality method, we will present a few illustrative applications of it. First, in Section 9.2, a pair of constant coefficient situations coming from [9, 26] for which comparison holds on all bounded domains and the needed (ZMP) for the $\widetilde{\mathcal{M}}$-subharmonics is accomplished by characterizing them. Then, in Section 9.3, we will discuss strict comparison for semiconvex functions in the variable coefficient setting and for semicontinuous functions in the constant coefficient setting.
@@ 9.2. Comparison with constant coefficients and sufficient monotonicity
We will present two basic examples of classes of $\mathcal{M}$-monotone subequations for which comparison holds on all bounded domains. The classes are determined by the two most elementary monotonicity cone subequations $\mathcal{M}$, which have interesting characterizations of the dual subharmonics ~$\widetilde{\mathcal{M}}(X)$, as known from [9, 26]. These characterizations easily lead to a direct proof that the $\widetilde{\mathcal{M}}$-subharmonics satisfy (ZMP) on all bounded domains, and hence the comparison principle in all $\mathcal{M}$-monotone potential theories. The two classes discussed here can be included in a much more general result that will be recalled in Remark 9.16 below.
@@
We have used Remark 5.15 to identify a constant coefficient subequation $\mathcal{F}\subset\mathcal{J}^{2}({\mathbb{R}}^{n})$ with a subset ×${\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$. Notice that such an $\mathcal{F}$ is indeed a subequation because the negativity property (N) is automatic and the topological property (T) follows from the $\mathcal{P}$-monotonicity, which implies the monotonicity property +(×$\mathcal{F}+({\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{P})\subset\mathcal{F}$, where the monotonicity +(×$\mathcal{F}+({\mathbb{R}}\times{\mathbb{R}}^{n}\times\{0\})\subset\mathcal{F}$ is an elegant way to say that $\mathcal{F}$ is of pure second order. In addition, since the only component of a jet that is actually relevant to define a subsolution is the one living in $\mathcal{S}(n)$, we will often identify $\mathcal{F}$ with $\mathcal{A}\subset\mathcal{S}(n)$ that satisfies the positivity condition (9.2). Hence, we can say that a pure second-order constant coefficient subequation is a closed subset $\mathcal{A}\subset\mathcal{S}(n)$ which is $\mathcal{P}$-monotone.
The most basic constant coefficient pure second order subequations can be defined in terms of the ordered eigenvalues
$\lambda_{1}(A)\leqslant\cdots\leqslant\lambda_{n}(A)$
of $A\in\mathcal{S}(n)$. Of particular importance are the convexity subequation
$\mathcal{P}\vcentcolon=\{A\in\mathcal{S}(n):\ A\geqslant 0\}=\{A\in\mathcal{S}(n):\ \lambda_{1}(A)\geqslant 0\},$
(see (6.4) and Example 5.17) and its dual subequation
$\widetilde{\mathcal{P}}\vcentcolon=\{A\in\mathcal{S}(n):\ -A\not\in\operatorname{int}\,\mathcal{P}\}=\{A\in\mathcal{S}(n):\ \lambda_{n}(A)\geqslant 0\},$
which is known as and the subaffine subequation, for reasons which we now review.
@@
$w\leqslant a\ \text{on $\partial\Omega$}\implies w\leqslant a\ \text{on $\Omega$}.$ (9.3)
@@
We have denoted the space of all affine functions by $\mathrm{Aff}$, without specifying their domain because it is known that any $a$ which is affine on an open subset $\Omega$ has a unique extension to the whole space ${\mathbb{R}}^{n}$ (see, e.g., [51]).
The following characterizations of $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ subharmonics are the content of [26, Proposition 4.5] (see also [9, Theorem 10.7]).
Proposition 9.5.
Let $X\subset{\mathbb{R}}^{n}$ be open and let $u\in\operatorname{USC}(X)$. Then one has the following:
  1. (a)
    $u\in\mathcal{P}(X)$ $\iff$ $u$ is locally convex (away from the connected components on which $u\equiv-\infty$);
  2. (b)
    $u\in\widetilde{\mathcal{P}}(X)$ $\iff$ $u$ is subaffine on $X$ (that is, $\widetilde{\mathcal{P}}(X)=\mathrm{SA}(X)$).
Notice that, for real-valued $u$, Proposition 9.5(a) yields
$u\ \text{locally convex in $X$}\iff u\in\mathcal{P}(X),$
which is the characterization of locally convex functions as those functions having non-negative Hessian in the viscosity sense, as we anticipated the discussion preceding Proposition 1.28.
[…]
second, denoting by $\overline{\mathrm{C}}\mathrm{onvex}(X)$ the set of upper semicontinuous functions which are locally either convex or identically $\infty$- ∞,
$u\in\overline{\mathrm{C}}\mathrm{onvex}(X)\iff u+v\in\mathrm{SA}(X)\ \ \forall\,v\in\mathrm{SA}(X).$
These are the content of [26, Propositions 2.5 and 2.6], respectively. Moreover, they are, in the end, consequences of the jet-addition formula
$\mathcal{P}+\widetilde{\mathcal{P}}\subset\widetilde{\mathcal{P}},$
which comes from the fact that $\mathcal{P}$ is a monotonicity cone for itself and hence also for the dual $\widetilde{\mathcal{P}}$. More precisely, they represent consequences of the most elementary example of the Subaffine Theorem 9.6 below, in the case where $\mathcal{F}=\mathcal{P}$.
Indeed, note that $\mathcal{P}$ is a monotonicity cone subequation for all pure second-order constant coefficient subequations (and it is minimal in this sense), hence a straightforward application of the Subharmonic Addition Theorem (Corollary 8.9) yields the following result (while for a proof that uses Słodkowski’s Largest Eigenvalue Theorem and semiconvex approximation, the reader can instead refer to [26]).
Theorem 9.6 (Subaffine Theorem; [26, Theorem 6.5]).
Let $\mathcal{F}$ be a pure second-order constant coefficient subequation. Then, for any open set $X\subset{\mathbb{R}}^{n}$,
@@
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded domain and $w\in\operatorname{USC}(\overline{\Omega})\cap\mathrm{SA}(\Omega)$. Then (9.3) holds for all $a\in\mathrm{Aff}$.
@@
Assume without loss of generality that $a=0$. We consider an exhaustion of $\Omega$ by an increasing sequence of compact sets $\{K_{j}\}_{j\in\mathbb{N}}$ and set $U_{\delta}\vcentcolon=\{x\in\overline{\Omega}:\,w(x)<\sup_{\partial\Omega}w+\delta\}$, for any $\delta>0$. Since $w\in\operatorname{USC}(\overline{\Omega})$, we know that $U_{\delta}$ is an open neighborhood of $\partial\Omega$ in $\overline{\Omega}$. Therefore $\partial K_{j}\subset U_{\delta}$ for all $j$ sufficiently large and, since $w$ is subaffine, $\sup_{K_{j}}\leqslant\sup_{\partial\Omega}w+\delta$. This proves that $\sup_{\Omega}w\leqslant\sup_{\partial\Omega}w+\delta$; since $\delta$ is arbitrary, $w\leqslant 0$ on $\partial\Omega$ and $w$ is continuous, we conclude that $w\leqslant 0$ in $\overline{\Omega}$, which is the desired conclusion. ∎
Theorem 9.8 (Comparison for pure second-order constant coefficient subequations).
@@
$\mathcal{R}+\mathcal{N}\subset\mathcal{R}\quad\text{and}\quad\mathcal{A}+\mathcal{P}\subset\mathcal{A},$
@@
Just as in Remark 9.2, we have identified the constant coefficient $\mathcal{F}$ with a subset of $\mathcal{J}^{2}$ and it is indeed a subequation, since the remaining topological property (T) again follows by the $(\mathcal{N}\times{\mathbb{R}}^{n}\times\mathcal{P})$-monotonicity of $\mathcal{F}$. Moreover, one can again reduce the subequation by suppressing the “silent” gradient variable, which then allows one to further identify $\mathcal{F}$ with a subset of ×${\mathbb{R}}\times\mathcal{S}(n)$. Hence, one can say that a gradient-free constant coefficient subequation is a closed subset ×$\mathcal{F}\subset{\mathbb{R}}\times\mathcal{S}(n)$ which is $\mathcal{Q}$-monotone where
$\mathcal{Q}\vcentcolon=\mathcal{N}\times\mathcal{P}\subset\mathcal{{\mathbb{R}}}\times\mathcal{S}(n).$ (9.4)
@@
$\widetilde{\mathcal{Q}}=\{(r,A)\in{\mathbb{R}}\times\mathcal{S}(n):\ r\leqslant 0\ \text{or}\ A\geqslant 0\}.$
Also in this case the name comes from a characterization of dual subharmonics: by [9, Theorem 10.7], an upper semicontinuous function on $X$ is $\widetilde{\mathcal{Q}}$-subharmonic if and only if it is subaffine-plus on $X$; that is,
@@
where $\mathrm{SA}^{+}(X)$ is the set of all subaffine-plus functions on $X$, which are defined as follows.
@@
the set of affine-plus functions on $K$.
@@
A function $w\in\operatorname{USC}(X)$ is called subaffine-plus on $X$, and we write $w\in\mathrm{SA}^{+}(X)$, if (9.3) holds for every open subset $\Omega\Subset X$ and each $a\in\mathrm{Aff}^{+}(\overline{\Omega})$.
Remark 9.13.
As shown in [9, Theorem 10.7], one has the following interesting characterization of the subaffine-plus functions on open subsets $X$ of ${\mathbb{R}}^{n}$:
$\mathrm{SA}^{+}(X)=\{u\in\operatorname{USC}(X):\ u^{+}\in\mathrm{SA}(X)\},$
so that the $\widetilde{\mathcal{Q}}$-subharmonics on $X$ are those upper semicontinuous functions whose positive part is subaffine on $X$; note that this is equivalent to
$\widetilde{\mathcal{Q}}(X)=\{u\in\operatorname{USC}(X):\ u^{+}\in\widetilde{\mathcal{P}}(X)\}.$
@@
Theorem 9.14 (Subaffine-Plus Theorem; [9, Theorem 10.8]).